tangents and normals find the equation of the tangent to the curve x^2 + 3y – 3 =0 which is parallel to the line y=4x – 5.

Grade:12

1 Answers

Jitender Singh IIT Delhi
9 years ago
Ans:
Hello Student,
Please find answer to your question below

3y = x^2 - 3
Slope of tangent: y’ = 4
3y' = 2x
3.4 = 2x
x = 6
3y = 6^2 - 3
y = 11
Equation of tangent:
(y-11) = 4(x-6)